OpenStudy (anonymous):

How many mL of C16H34 are needed to react with 0.05 moles of O2? The density of C16H34 is 0.80g/mL.

OpenStudy (anonymous):

C16H34 + O2 v = ? mL n = 0.05 moles d = 0.8 g/mL Since it's a 1 to 1 ratio between the two reactants, you can use the number of moles that you have for O2 and use it for for C16H34. Also, you can use C16H34's molar mass, 226.34 I believe, to find the mass of C16H34. \[m = M \times n\] With the mass and density (which was given), you can find volume using the formula: \[v =\frac{ m }{ d }\] Ta-da. :- )

OpenStudy (anonymous):

Thank you!